# SPECIAL RELATIVITY

### v11

@@ 2.9 Commonly Used Method of Coupling Fields and Particles
×$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}~{},$ (7)
×$\vec{\nabla}\times\vec{B}=\frac{4\pi}{c}\vec{j}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}~{}.$ (8)
@@ 3.1 Choosing a Coordinate System in an Inertial Frame
There are multiple methods to synchronize distant clocks [^8] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity between distant events is a matter of convention, as it can be legitimately defined in different ways within any given inertial reference frame. Moeller [13], for example, states: ”All methods for the regulation of clocks meet with the same fundamental difficulty. The concept of simultaneity between two events in different places has no exact objective meaning at all since we cannot specify any experimental method by which this simultaneity could be ascertained. The same is therefore true also for the concept of velocity.”. The choice of a synchronization convention is ultimately a matter of selecting a coordinate system within an inertial frame of reference in Minkowski space-time.
@@ 3.2 The Inertial Frame Where a Light Source is at Rest
Our goal is to provide a practical and operational procedure for this assignment. The most natural synchronization method involves first bringing all ideal clocks to the same spatial location, where they can be synchronized. Once synchronized, the clocks are then slowly transported to their designated positions—a process known as slow clock transport  [^9] Eddington, in a text first published in 1923, was apparently the first to describe a synchronization procedure based on the slow transport of clocks [18]. Further details can be found in the review [8] (see also [19])..
@@
Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest  [^10] Many textbook authors still attribute a reality status to the one-way speed of light. To quote Hrasko [20]: ”It is sometimes claimed that Einstein’s synchronization of distant clocks $A$ and $B$ is circular. … This argument, however, is fallacious. It is true that a one-way measurement of light velocity can be performed only if clocks at the endpoints are correctly synchronized. However, since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them symmetrically from a common site. … As we see, the described thought experiment is capable of proving the constancy of the speed of light if it is true or disproving it if it is false. It therefore provides a solid logical foundation for Einstein’s synchronization prescription.” This reasoning, however, is incorrect. Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest.. In other words, suppose we have two sets of synchronized clocks spaced along the $x$-axis. Suppose that one set of clocks is synchronized by using the slow clock transport procedure and the other by light signals. If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared. This is because in our case of interest when the light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Applying either synchronization method in the rest frame establishes a Lorentz coordinate system. In this system, the metric of the light source takes the Minkowski form. We now turn to the topic of electromagnetic waves. In free space, the electric field $\vec{E}$ of an electromagnetic wave emitted by a stationary dipole source satisfies the wave equation: $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. This result is closely related to the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
@@ 3.3 Motion of a Light Source Relative to an Inertial Frame
This metric describes the electrodynamics of the moving light source from the viewpoint of the inertial observer measurements. From the structure of the equation, it is evident that the measurement of the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a measuring device at rest in the inertial frame—yields the same result as that obtained at the spatial position $x-vt$ at time $t=0$.
@@
The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$. Here we obtained the solution for waves that move in the $x$ direction by supposing that the field does not depend on $y$ and $z$. The first term represents a wave traveling forward in the positive $x$ direction, and the second term represents a wave traveling backward in the negative direction.
We conclude that the speed of light emitted by a moving source, as measured in the laboratory frame $(t,x,y,z)$, depends on the relative velocity $v$ between the source and the observer. In other words, in this scenario, the speed of light follows the Galilean law of velocity addition. Specifically, the coordinate velocity of light along the $x$-axis is given by $dx/dt=c+v$ in the positive direction and $dx/dt=-c+v$ in the negative direction. This apparent deviation from the electrodynamics constant $c$ arises because the clocks are synchronized according to an absolute time convention. Once such a convention is chosen, it must be consistently applied to both mechanics and electrodynamics. Under this absolute time coordinatization, the velocity of light transforms analogously to that of a particle, as expected.
@@ 3.4 Discussion
3.5 A Clock Resynchronization Procedure
After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration. The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation. To get around this difficulty, we observe that simplification is always possible. The trick needed here is to further make a change of the time variable. Suppose that we describe the effect of radiation by working only up to the first order in $v/c$. In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$. After this time shift we obtain the usual wave equation.
@@ 3.5 A Clock Resynchronization Procedure
Consider two light sources labeled ”1” and ”2”. Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$. The question now arises of how to assign a time coordinate to the lab reference frame. We have two choices: an absolute time coordinate or a Lorentz time coordinate. The most natural choice, in terms of connecting to laboratory reality, is absolute time synchronization. In this framework, simultaneity is absolute, meaning that a single set of synchronized clocks is sufficient for both sources in the lab frame. However, Maxwell’s equations are not form-invariant under Galilean transformations, meaning their mathematical form differs in the lab frame. Using absolute time synchronization leads to more complex field equations, which also vary for each source. To introduce Lorentz coordinates, we must adopt a different approach. The only viable method in this scenario is to assign an individual coordinate system—i.e., an independent set of synchronized clocks—to each source.
Now, consider a second secondary source, such as a tangentially moving mirror, interacting with a light beam emitted by
a stationary source. A key challenge in special relativity is that both sources cannot share a common Lorentz frame within a single inertial system. The electrodynamic equations must be identical for both the incident and reflected waves. Under Einstein’s synchronization, the time for the moving mirror can be determined by introducing a time shift of $xv/c^{2}$ to the first order of approximation. In this synchronization scheme, reflection is described using Maxwell’s equations. This time shift effectively rotates the radiation wavefront by an angle $v/c$ for both sources. Now, consider the electrodynamics of the stationary source. In absolute-time coordinatization, an observer describes the light beam from the stationary emitter using the diagonal wave equation. However, the time shift introduces a new crossed term in the wave equation, leading to a non-diagonal form. Further details on this issue are provided in the next chapter (see Section 4.6 for further details).
@@ 4 Aberration of Light in an Inertial Frame of Reference
The effect of light aberration in an inertial frame of reference is typically understood as a change in the direction of light propagation attributed to moving light sources. However, even at the first order in $v/c$, describing this phenomenon rigorously is far from straightforward. Many textbooks contain incorrect results, reflecting the complexity of the topic. To minimize mathematical complexity, we will restrict our discussion to first-order effects. This chapter focuses on the physical influence of optical instruments on the measurement of light aberration. First, we will demonstrate that when a finite-aperture mirror moves parallel to itself and a plane wave of light falls normally onto it, the reflected light exhibits aberration—manifested as a deviation in energy transport. The next problem, of greater practical significance, involves the complementary case of transmission through an aperture in a moving opaque screen, as described by Babinet’s principle. A particularly relevant scenario is the transmission of light through the moving end of a telescope barrel, which we will examine in detail.
@@ 4.1 The ”Plane Wave” Emitter
We will analyze the effect in a simple case to build an intuitive understanding. The explanation of aberration is based on modeling a single plane-wave emitter. As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$. We take the elementary sources to start radiating waves simultaneously with respect to the lab reference frame where the plane $P$ is at rest. Therefore we have a plane full of sources, oscillating together, with their motion in the plane and all having the same amplitude and phase. Let us assume that the elementary sources are oscillating at frequency $\omega$. By reducing the spacing between adjacent sources to a value much smaller than the radiation wavelength $\lambda=2\pi c/\omega$, we can approximate this two-dimensional arrangement as an ideal plane-wave emitter.
The concept of an (infinitely extended) plane wave is widely used in physics due to its well-behaved analytical properties as a solution to Maxwell’s equations. However, it is not physically realizable, as such a wave would carry infinite energy. In any real-world scenario, one must consider a finite source aperture. It is implicitly assumed that the detector measuring the radiation direction is sensitive to energy propagation and that its aperture is sufficiently large compared to the radiation beam size. Ultimately, what is commonly referred to as aberration is, in reality, a deviation in the direction of energy transport.
4.2 A Moving Emitter: Single Frame Description
Let us consider the case of an emitter lying in the $(x-y)$ plane of an inertial laboratory frame, which is accelerated from rest to a velocity $v$ along the $x$-axis. Suppose an observer at rest in this inertial frame measures the direction of energy transport.
There are two possible approaches to analyzing the aberration of light emitted by a single moving source. The first is the covariant approach, commonly used in the literature, which explains the aberration effect by applying a Lorentz boost to determine how the direction of a light beam depends on the velocity of the source relative to the inertial frame. The second is a non-covariant approach, which employs a ”single-frame” description without explicit reference to Lorentz transformations.
@@ 4.2 A Moving Emitter: Single Frame Description
Let us examine the electrodynamics of a moving source. To explain the phenomenon of radiation in our case of interest, we apply an (active) Galilean boost to describe the uniform translational motion of the source in the inertial lab frame. Since Maxwell’s equations are not invariant under Galilean transformations, their form is not preserved. The additional terms introduced in the field equations as a result of the Galilean transformation lead to the prediction of the Doppler effect.
The current approach to the moving emitter problem employs the Fourier transform method. In the analysis of linear systems, it is often useful to break down a complex input into simpler components, determine the system’s response to each elementary function, and then superimpose these responses to obtain the total outcome. Fourier analysis serves as a fundamental tool for this decomposition  [^11] For a broader discussion on Fourier transform methods in spatial filtering theory or Abbe diffraction theory, we recommend referring to [22]..
[…]
An emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the current density of the elementary dipole sources varies according to the law $j_{pol}=g(K_{\perp})\sin{({K}_{\perp}x)}$. From this, we conclude that the active medium of the emmiter is sinusoidally space-modulated.
@@
The wavevector $\vec{k}$ is determined by the initial conditions. Specifically, we set $k_{x}=K_{\perp}$, where $K_{\perp}$ corresponds to the wavenumber of the sinusoidally space-modulated dipole density, and $k_{z}^{2}=\omega_{i}^{2}/c^{2}$, with $\omega_{i}$ denoting the frequency of the emitted radiation prior to acceleration.
Given these initial conditions, we treat $\vec{k}$ as the independent variable and express the frequency as a function of $k_{x}$: $\omega=\omega_{i}+\Delta\omega(k_{x})$, where $\Delta\omega(k_{x})$ represents the frequency shift due to motion. Substituting this relation into the dispersion equation yields a condition linking the wavenumber and the frequency shift: $\Delta\omega=K_{\perp}v$.
[…]
An alternative physical perspective on the Doppler effect for a light beam emitted by a moving source leads to the same conclusion as the one presented above. The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system. This invariance holds regardless of the specific coordinate transformation used to describe the change in the reference system. Therefore, the phase $\vec{k}\cdot\vec{r}-\omega t$ must be invariant of the Galilean transformation. Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$. Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\Delta\omega=K_{\perp}v$. This frequency shift matches the result obtained directly from the dispersion relation, as expected.

### v12

@@ 2.9 Commonly Used Method of Coupling Fields and Particles
$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}~,$ (7)
$\vec{\nabla}\times\vec{B}=\frac{4\pi}{c}\vec{j}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}~.$ (8)
@@ 3.1 Choosing a Coordinate System in an Inertial Frame
There are multiple methods to synchronize distant clocks. [^8] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity between distant events is a matter of convention, as it can be legitimately defined in different ways within any given inertial reference frame. Moeller [13], for example, states: ”All methods for the regulation of clocks meet with the same fundamental difficulty. The concept of simultaneity between two events in different places has no exact objective meaning at all since we cannot specify any experimental method by which this simultaneity could be ascertained. The same is therefore true also for the concept of velocity.” The choice of a synchronization convention is ultimately a matter of selecting a coordinate system within an inertial frame of reference in Minkowski space-time.
@@ 3.2 The Inertial Frame Where a Light Source is at Rest
Our goal is to provide a practical and operational procedure for this assignment. The most natural synchronization method involves first bringing all ideal clocks to the same spatial location, where they can be synchronized. Once synchronized, the clocks are then slowly transported to their designated positions—a process known as slow clock transport.  [^9] Eddington, in a text first published in 1923, was apparently the first to describe a synchronization procedure based on the slow transport of clocks [18]. Further details can be found in the review [8] (see also [19]).
@@
Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest.  [^10] Many textbook authors still attribute a reality status to the one-way speed of light. To quote Hrasko [20]: ”It is sometimes claimed that Einstein’s synchronization of distant clocks $A$ and $B$ is circular. … This argument, however, is fallacious. It is true that a one-way measurement of light velocity can be performed only if clocks at the endpoints are correctly synchronized. However, since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them symmetrically from a common site. … As we see, the described thought experiment is capable of proving the constancy of the speed of light if it is true or disproving it if it is false. It therefore provides a solid logical foundation for Einstein’s synchronization prescription.” This reasoning, however, is incorrect. Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest. In other words, suppose we have two sets of synchronized clocks spaced along the $x$-axis. Suppose that one set of clocks is synchronized by using the slow clock transport procedure and the other by light signals. If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared. This is because in our case of interest when the light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Applying either synchronization method in the rest frame establishes a Lorentz coordinate system. In this system, the metric of the light source takes the Minkowski form. We now turn to the topic of electromagnetic waves. In free space, the electric field $\vec{E}$ of an electromagnetic wave emitted by a stationary dipole source satisfies the wave equation: $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. This result is closely related to the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
@@ 3.3 Motion of a Light Source Relative to an Inertial Frame
This metric describes the electrodynamics of the moving light source from the viewpoint of the inertial observer measurements. From the structure of the equation, it is evident that the measurement of the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a measuring device at rest in the inertial frame—yields the same result as that obtained at the spatial position xvtx-vt at time $t=0$.
@@
The solution of this equation F[x(c+v)t]+G[x+(c+v)t]F[x-(c+v)t]+G[x+(-c+v)t] is the sum of two arbitrary functions, one of argument x(c+v)tx-(c+v)t and the other of argument x+(c+v)tx+(-c+v)t. Here we obtained the solution for waves that move in the $x$ direction by supposing that the field does not depend on $y$ and $z$. The first term represents a wave traveling forward in the positive $x$ direction, and the second term represents a wave traveling backward in the negative direction.
We conclude that the speed of light emitted by a moving source, as measured in the laboratory frame $(t,x,y,z)$, depends on the relative velocity $v$ between the source and the observer. In other words, in this scenario, the speed of light follows the Galilean law of velocity addition. Specifically, the coordinate velocity of light along the $x$-axis is given by $dx/dt=c+v$ in the positive direction and dx/dt=c+vdx/dt=-c+v in the negative direction. This apparent deviation from the electrodynamics constant $c$ arises because the clocks are synchronized according to an absolute time convention. Once such a convention is chosen, it must be consistently applied to both mechanics and electrodynamics. Under this absolute time coordinatization, the velocity of light transforms analogously to that of a particle, as expected.
@@ 3.4 Discussion
At this point, a reasonable question arises: Why address this topic now? Why not wait until Chapter 5? The answer is that the subject is subtle and complex, and the best way to understand it is through gradual exposure. The first step is to develop an intuitive sense of what should occur in various situations. Later, with a more solid grasp of relativistic effects, we will revisit the topic and refine our understanding with greater precision.
3.5 A Clock Re-Synchronization Procedure
Most existing studies assume that an observer at rest and a moving source share the same Minkowski metric, thereby neglecting the need for any clock re-synchronization procedure. However, this standard approach is flawed, as it fails to account for the principle of equivalence between active and passive boosts within a single inertial frame.
Our analysis shows that after applying a Galilean boost under an absolute time coordinatization, the homogeneous wave equation for the field in the lab frame appears nearly—but not exactly—in its standard form (valid in the absence of acceleration). The key deviation lies in the presence of a mixed derivative term, $\partial^{2}/\partial t,\partial x$, which complicates the equation’s solution.
To resolve this, we note that a simplification is always possible by introducing a suitable change of the time variable. If we limit our analysis to first-order terms in $v/c$, a time transformation of the form $t\to t+xv/c^{2}$ can be applied. This time shift effectively eliminates the mixed derivative term, reducing the equation to the standard wave equation form.
@@ 3.5 A Clock Re-Synchronization Procedure
Consider two independent light sources labeled ”1” and ”2”. Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$. The question now arises of how to assign a time coordinate to the lab reference frame. We have two choices: an absolute time coordinate or a Lorentz time coordinate. The most natural choice, in terms of connecting to laboratory reality, is absolute time synchronization. In this framework, simultaneity is absolute, meaning that a single set of synchronized clocks is sufficient for both sources in the lab frame. However, Maxwell’s equations are not form-invariant under Galilean transformations, meaning their mathematical form differs in the lab frame. Using absolute time synchronization leads to more complex field equations, which also vary for each source. To introduce Lorentz coordinates, we must adopt a different approach. The only viable method in this scenario is to assign an individual coordinate system—i.e., an independent set of synchronized clocks—to each source.
The situation becomes more complex when a secondary source interacts with light emitted by a primary source. A fundamental challenge in special relativity is that both sources cannot, in general, be described using a single shared Lorentz frame within a common inertial reference system. This issue, and its implications for electrodynamics, is discussed in more detail in the next chapter.
@@ 4 Aberration of Light in an Inertial Frame of Reference
In this chapter, we explore several practical applications of the concepts introduced earlier, using the spacetime metric given in Eq. (11). Our first topic addresses the phenomenon of light aberration caused by a light source moving perpendicular to the direction of radiation.
Light aberration in an inertial frame is generally understood as an apparent shift in the direction of light propagation due to the motion of the source. While often treated as a first-order effect in $v/c$, a rigorous description is not trivial—even at this level of approximation. Many standard textbooks contain inaccurate or oversimplified accounts, underscoring the subtlety of the underlying physics. To keep the analysis manageable, we will limit our discussion to first-order effects in $v/c$.
Our focus will be on the physical influence of optical components on the observed aberration of light. We begin by examining the case of a finite-aperture mirror moving parallel to its surface, with an incident plane wave striking it normally. We will show that the reflected light exhibits aberration, observable as a deviation in the direction of energy transport.
We then turn to a more practically significant scenario: the transmission of light through an aperture in a moving opaque screen, invoking Babinet’s principle. A particularly important application is the passage of light through the moving entrance of a telescope barrel—a configuration we will analyze in detail.
@@ 4.1 The ”Plane Wave” Emitter
To build intuitive insight into the phenomenon of aberration, we begin by analyzing a simple, idealized case: a single ”plane-wave” emitter.
We model this emitter as a two-dimensional array of identical, coherent elementary sources (dipoles), uniformly distributed over a fixed (xxyy) plane. These sources are assumed to begin radiating simultaneously in the lab frame, where the plane is at rest. All dipoles oscillate with the same frequency $\omega$, amplitude, and phase, and their motion is confined to the plane.
By decreasing the spacing between adjacent sources to a value much smaller than the radiation wavelength, $\lambda=2\pi c/\omega$, this arrangement effectively approximates an ideal plane-wave emitter. The resulting wavefronts are planar, and the emitted radiation resembles an ideal plane wave.
While the concept of an (infinitely extended) plane wave is a useful mathematical abstraction—owing to its simplicity and exact correspondence with solutions to Maxwell’s equations—it is not physically realizable, as such a wave would carry infinite energy. In any real scenario, the emitter has a finite aperture.
It is also assumed that any detector used to measure the radiation direction is sensitive to energy flow (rather than to phase fronts, for instance) and has an aperture large enough compared to the beam size to capture the overall direction of energy propagation. What is commonly described as aberration, in this context, is in fact a shift in the direction of energy transport.
4.2 A Moving Emitter: Galilean Transformations-Based Explanation
Let us consider the case of an emitter lying in the (xy)(x-y) plane of an inertial laboratory frame, which is accelerated from rest to a velocity $v$ along the $x$-axis. Suppose an observer at rest in this inertial frame measures the direction of energy transport.
There are two possible approaches to analyzing the aberration of light emitted by a single moving source. The first is the covariant approach, commonly used in the literature, which explains the aberration effect by applying a Lorentz boost to determine how the direction of a light beam depends on the velocity of the source relative to the inertial frame. The second is a non-covariant approach, which employs an old kinematics description without explicit reference to Lorentz transformations.
@@ 4.2 A Moving Emitter: Galilean Transformations-Based Explanation
The current approach to the moving emitter problem employs the Fourier transform method. In the analysis of linear systems, it is often useful to break down a complex input into simpler components, determine the system’s response to each elementary function, and then superimpose these responses to obtain the total outcome. Fourier analysis serves as a fundamental tool for this decomposition.  [^11] For a broader discussion on Fourier transform methods in spatial filtering theory or Abbe diffraction theory, we recommend referring to [22].
Consider the inverse transform relationship $g(x)=\int_{-\infty}^{\infty}G(K)\exp(iKx)dK$ which express the profile function in terms of its wavenumber spectrum. We may regard this expression as a decomposition of the function $g(x)$ into a linear combination (in our case into an integral) of elementary functions, each with specific form $\exp(iKx)$. From this, it is clear that the number $G(K)$ is simply a weighting factor that must be applied to the elementary function of wavenumber $K$ to synthesize the desired $g(x)$.
An emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the current density of the elementary dipole sources varies according to the law $j_{pol}=g(K_{\perp})\sin{({K}_{\perp}x)}$. From this, we conclude that the active medium of the emitter is sinusoidally space-modulated.
Let us examine the electrodynamics of a moving source. To explain the phenomenon of radiation in our case of interest, we apply an (active) Galilean boost to describe the uniform translational motion of the source in the inertial lab frame. According to equivalence of passive and active boosts, Maxwell’s equations remains valid in the comoving frame. The measurement of the electromagnetic field configuration - expressed in terms of coordinates $x$ and $t$ of a measuring device at rest - yields the same result that obtained at the spatial position xvtx-vt at time $t=0$. Applying a Galilean boost, we transform the spatial coordinate as $x\to x-vt$, while keeping time unchanged. Since Maxwell’s equations are not invariant under Galilean transformations, their form is not preserved.
@@
The wavevector $\vec{k}$ is determined by the initial conditions before the acceleration and subsequent Galilean transformation. Specifically, we set $k^{\prime}_{x}=K_{\perp}$, where $K_{\perp}$ corresponds to the wavenumber of the sinusoidally space-modulated dipole density, and $(k_{z}^{\prime})^{2}=\omega_{i}^{2}/c^{2}-K_{\perp}^{2}$, with $\omega_{i}$ denoting the frequency of the emitted radiation prior to acceleration.
The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system (when the field is zero, everyone measures the field as zero). This invariance holds regardless of the specific coordinate transformation used to describe the change in the reference system. Therefore, the phase $\vec{k}\cdot\vec{r}-\omega t$ must be invariant of the Galilean transformation. Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$. Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\omega=\omega_{i}+K_{\perp}v$. This transformation completes with the invariance of wavenumber vector $k_{x}=K_{\perp}$, $k_{z}^{2}=\omega_{i}^{2}/c^{2}-K_{\perp}^{2}$. Substituting these into dispersion relation confirms that it remains satisfied, as expected.